A rectangle is measured to be 6.4 ± 0.2 cm by 8.3 ± 0.2 cm. Calculate the uncertainty in its perimeter.
Answer #1

lit e = 6.48.0.2 cm, bi 8.31 0.2cm o Perimeter P = zle+b). > Po 2(6.4 +8.3) a PE 29.4 cm A P= 21+ 2b. :: lop= 216 2+ zhbb sis